OpenStudy (anonymous):

suppose \frac{ 1 }{ a^{2} } - \frac{ 1 }{ b ^{2} } = \frac{ 1 }{ c } , ab=sqrt{c} , then the value of the average of 'a' and 'b' is?????

OpenStudy (anonymous):

?\[suppose \frac{ 1 }{ a^{2} } - \frac{ 1 }{ b ^{2} } = \frac{ 1 }{ c } , ab=\sqrt{c} , then the value of the average of 'a' and 'b' is?????\]